OpenStudy (anonymous):

What is the remainder when (2x4 – 3x3 + x2 – 3x – 3) ÷ (x – 2)

sam (.sam.):

Start with 

sam (.sam.):

Or you could use f(x)=2x4 – 3x3 + x2 – 3x – 3, When (x-2)=0, x=2 Then f(2)=2(2)4 – 3(2)3 + (2)2 – 3(2) – 3=?

OpenStudy (anonymous):

-9

sam (.sam.):

Careful \[2(2)^4 – 3(2)^3 + (2)^2 – 3(2) – 3=?\]

OpenStudy (anonymous):

@.Sam. 6

sam (.sam.):

Nope

OpenStudy (anonymous):

@.Sam. well idk

sam (.sam.):

\[2^4=2\times 2 \times 2 \times 2=16\] \[2^3=2 \times 2 \times 2 = 8\] \[2^2=2 \times 2 = 4\] ----------------------------------------------- So, \[2(2)^4–3(2)^3+(2)^2–3(2)–3 \\ 2(16)-3(8)+4-6-3 \\ =3\]